OpenStudy (strawberry17):

how would you write 0.01% as a decimal. Would it be 0.001 or 0.0001. And please explain. Thanks!

jimthompson5910 (jim_thompson5910):

x% = x/100 0.01% = 0.01/100 0.01% = 0.0001

OpenStudy (strawberry17):

okay thanks so much!

jimthompson5910 (jim_thompson5910):

np